Oero ting profit Margins 7. interesst9,000 tax - 9,000 osset tunoues Find ROE ASSETSSALES ASSETSSALES OPM Aug. EQ
Answer #1

Given

  • Operating profit, EBIT = 36,000
  • Sales
    • Operating margin = 0.07 = EBIT/Sales
    • 0.07 = 36,000/ Sales
    • Sales = 36,000 / 0.07 = 514,285.7
    • Sales = 514,286 (rounding to nearest integer)
  • Debt/Equity = 1.45
    • Capital = Debt+Equity = 1.45 + 1 = 2.45
    • Debt/Capital = 1.45/2.45 = 0.59
    • Equity/Capital = 1/2.45 = 0.41
  • Net Income
    • EBIT = 36,000
    • Less: Interest = 9,000
    • EBT = 36,000 - 9,000 = 27,000
    • Less: Tax = 9,000
    • EAT = Net income, NI = 18,000
  • Asset Turnover = 4
    • Asset turnover = Sales/Total Assets = 4
    • Sales = 514,286
    • Total Assets = Sales / 4 = 514,286/4 = 128,571
  • Shareholder's Equity
    • Total Assets = 128,571
    • Shareholder's equity = Total assets / 2.45 = 128,571/2.45 = 52,478

ROE

a)ROE = (TA/Sh.Equity) * (Sales/Assets) * (OPM/Sales) * (Net Income/OPM)

  • ROE = (128,571/52,478) * (514,286/128,571) * (0.07/514,286) * (18,000/0.07)
  • ROE = 34.30%

b)NI/Avg Equity = 18,000/52,478 = 0.3430 = 34.30%

c)ROE = (Avg Assets/Avg Equity) * (Sales/ Assets) * (Net Income/ Sales)

  • ROE = (128,571/52,478) * (514,286/128,571) * (18,000/514,286)
  • ROE = 2.45 * 4 * 0.04 = 0.343 = 34.30%